OpenStudy (haleyelizabeth2017):

If α and β are the measures of two first quadrant angles and sin α =4/5 and sin β=5/13, find sin (α+β)

OpenStudy (mathstudent55):

Are you familiar with the trig addition identities? This is the one you need. \(\sin (\alpha + \beta) = \sin \alpha \cos \beta + \sin \beta \cos \alpha \) First, you need to find the cosines of those two angles.

OpenStudy (haleyelizabeth2017):

Okay

OpenStudy (haleyelizabeth2017):

cos α is 3/5 right?

OpenStudy (haleyelizabeth2017):

and cos β is 12/13

OpenStudy (mathstudent55):